OpenStudy (anonymous):

what is the proportion of graph paper 8.5 by 11 to final project of 170 by 220

OpenStudy (mokeira):

\[\frac{ 170 }{ 8.5 }=20\] \[\frac{ 220 }{ 11 }=20\]

OpenStudy (anonymous):

I got 1 so is the ratio 1 to 1?

OpenStudy (mokeira):

yeah :)

OpenStudy (anonymous):

I am creating a design on graph paper and have to state ratio of my design to the actual size of 170:220. How do I do that with a ratio of 1:1

OpenStudy (anonymous):

If my design is 8:1 how do I convert to actual size

OpenStudy (mokeira):

ratio is 1:20 where 20 is for actual

OpenStudy (anonymous):

how did u get 20

OpenStudy (anonymous):

so if my design on the graph paper is 2 x 2 do I set it up 1/20=2/x